(1/4 complete with respect to direct labor lete with respect to direct labor and Sewing Departments 19. A company uses a process cost accounting system. Its 120,000 units this period. Its beginning inventory consisted of 50,000 units FIFO and overhead). The Sewing Department started and finished ending inventory consists of 40,000 units (1/4 comp overhead). All direct materials are added at the beginning of the process. Under what are the equivalent units of production for the Sewing Department for direct materials and for direct labor and overhead, respectively? A. 210,000; 120,000 B. 210,000; 180,000 C. 167,500; 167,500 D. 160,000; 162,500 E. 160,000; 167,500 20. Total variable costs change proportionately with changes in output activity A. False B. False 21. Variable costs per unit increase proportionately with increases in output activity A. True B. False 22. A company manufactures and sells a product for $120 per unit. The companys fixed costs are $68,760, and its variable costs are $90 per unit. The companys break-even point in units is: A. 2,292 B. 573 C. 764 D. 327 E. 840 23. A company manufactures and sells a product for $91 per unit. The companys fixed costs are $859,716 and its variable costs are $25 per unit. The companys break-even point in units is: A. 7,412 B. 34,389 C. 9,448 D. 13,026 E. 66 24. A company manufactures and sells a product for $120 per unit. The companys fixed costs are $68,760, and its variable costs are $90 per unit. The companys break-even point in dollars is: A. $91,680 B. $68,760 C. $2,2921
Answer #1
19
Direct materials Direct labor and Overhead
To complete beginning inventory 0 37500 =50000*3/4
Units started and finished 120000 120000
Ending inventory 40000 10000 =40000*1/4
Equivalent units of Production 160000 167500
Option E is correct
20
True, total variable costs change proportionately with changes in output activity
21
False, variable costs per unit do not increase proportionately with changes in output activity. Variable costs per unit remain constant.
22
Break-even point in units = Fixed costs/Unit Contribution margin = 68760/(120-90)= 2292
Option A is correct
23
Break-even point in units = Fixed costs/Unit Contribution margin = 859716/(91-25)= 13026
Option D is correct
24
CM ratio = (120-90)/120 = 25%
Break-even point in units = Fixed costs/CM ratio = 68760/25%= $275040